# Biomedical open source software: Crucial packages and hidden heroes

### v4

@@ 1 Introduction
Besides this, there is another kind of invisibility of scientific software. The programs visible to the end user may rely on many other software packages (known in the software world as dependencies). While the end users may mention a package at the top of the dependency stack, they are likely not even aware of the full set of packages that are further below. These may be packages that the utilized program depends on directly (direct dependencies), or indirectly, where the direct dependencies in turn may depend on further packages. These latter packages thus become indirect, transitive dependencies of the user-facing program. This complex nature of the network of dependencies has a number of implications, including for security [Hatta2022Nebraska, 13] and computational reproducibility [14]. In particular, all the work undertaken to develop, maintain, test, and distribute the underlying software is not directly visible in the publication record itself, and is not included in derivatives such as citation networks and knowledge graphs.
The situation resembles the famous XKCD cartoon where “all modern infrastructure” critically depends on “a project some random person in Nebraska has been thanklessly maintaining since 2003” [1]. The word “thanklessly” is important in this context: being unknown, these critical pieces of software get much less recognition and credit than they deserve—and than the science needs. The absence of recognition may lead to dire consequences, for example, if, due to the lack of funding, the underlying libraries become vectors for malware attacks.
There have been proposals to assign credit to these packages using the dependency structure to calculate what was been termed transitive credit [15, 16] but so far, the problem has not yet been solved. To implement the proposed measure, software projects must start to publicize the packages they rely on in a way that enables recognition and citation, beyond the technical dependency already recorded in manifest files such as pyproject.toml, DESCRIPTION, cargo.toml, pom.xml, etc. One way for projects to do this is the inclusion of citation information for their own software outputs as well as for their direct dependencies, e.g., a citation file in the Citation File Format (CFF) [17]. Were such citation information available for the complete dependency stack of a program, transitive credit could be implemented by building weighted software citation networks [18].
@@
Following these ideas, we can put the software packages into a two-dimensional plane with the axes corresponding to the frequency of software mentions and a measure of network centrality (Figure 1). The majority of the packages will likely occupy the lower left corner of the plot, having a small number of mentions from the authors of scientific papers and limited centrality from usage in other software packages. Popular packages are used by many authors. We are interested in the “Nebraska” packages, not very visible, but of critical importance. Of course, one can think about highly visible and critical “Pasteur” packages, which are used both directly and as a foundation for other libraries.
@@ 3 Results
We found a dense core of popular packages that receive many mentions (e.g. ggplot2 in CRAN, tophat in PyPI and limma in Bioconductor), some of which have many dependencies themselves (e.g. ggplot2). We also found packages specific to certain communities (e.g. PRISMA in CRAN or pymol in PyPI).
@@
“Pasteur”: the highest sum of mentions and centrality percentile.
@@ 4 Discussion
The network analysis is limited by the quality of disambiguation and linking in the CZ Software Mentions dataset (see the discussion in [11]). Homonyms (different software packages with the same or close names) may change the network statistics, and preliminary research points to considerable potential for incorrect linking in the CZ Software Mentions dataset due to homonyms [7]. Notably, this affects the package velvet (see Figure 2a), which has been falsely linked to a Python package that provides signal processing and communications algorithms in Python [41]. Meanwhile, all of the mentions of “velvet” point to different versions of algorithms for de novo short-read assembly using de Bruijn graphs in genomics, as described in [42] and implemented mostly in C [40]. While this does not change its centrality and our interpretation, the linking error led to its mislabelling as a PyPI package for the visualization in Figure 2.
@@
Our approach is limited to the open-source packages available in one of the chosen ecosystems. This excludes proprietary software like Excel or GraphPad Prism, and open-source non-package software like Gephi [43] (a complied application). The number of missing packages may be estimated from Table 14 of [11], which estimates the coverage for GitHub at 64.39%, the coverage for CRAN at 8.36%, the coverage for PyPI at 5.86%, and the coverage for Bioconductor at 3.23%. GitHub coverage can be used as a proxy for the share of the open source software: while some packages—perhaps growing in number—are not hosted by GitHub, their fraction remains low [44]. If we assume that each package exists only in one ecosystem (there are packages present in several, but we neglect this effect), the combined coverage by CRAN, PyPI, and Bioconductor is 17.45%, i.e., between a quarter and one third of the open source packages. It is very likely that each of these non-package pieces of software do depend on packages in our focus ecosystems. For open-source non-package software, it may be feasible to identify source repository URLs and then identify packages depended on by the software. That would need to be done by directly inspecting the source code, possibly using Software Bill of Materials (SBOM) tools [Hatta2022Nebraska]. On the other hand, some non-package software does not have source code available (including proprietary GUI software such as SPSS, and cloud-based services). Nonetheless, these pieces of code may very well be based on packages in our focal ecosystem (especially if the latter use non-restrictive licenses such as BSD, MIT). In the long term, there is a possibility that the requirement to provide SBOMs for U.S. federal government purchasers111as per Executive Order 14028: Improving the Nation’s Cybersecurity from May 12, 2021 will give insight into the packages on which proprietary software relies [45].
@@
There are different software package repositories and ecosystems for Python and R with greatly overlapping dependency graphs. We used PyPI, CRAN and Bioconductor for this study. Alternative repositories that provide packages for Python and R, such as Conda channels (conda-forge, bioconda, etc.), are also worth investigating.

### v5

@@ 1 Introduction
Besides this, there is another kind of invisibility of scientific software. The programs visible to the end user may rely on many other software packages (known in the software world as dependencies). While the end users may mention a package at the top of the dependency stack, they are likely not even aware of the full set of packages that are further below. These may be packages that the user-facing program depends on directly (direct dependencies), or indirectly, where the direct dependencies in turn may depend on other packages. These latter packages thus become indirect, transitive dependencies of the user-facing program. This complex nature of the network of dependencies has a number of implications, including those for security [12, 13] and computational reproducibility [14]. In particular, much of the work undertaken to develop, maintain, test, and distribute the underlying software is not directly visible in the publication record itself, and is not included in derivatives such as citation networks and knowledge graphs.
The situation resembles the famous XKCD cartoon where “all modern infrastructure” critically depends on “a project some random person in Nebraska has been thanklessly maintaining since 2003” [15]. The word “thanklessly” is important in this context: being unknown, these critical pieces of software get much less recognition and credit than they deserve—and than the science needs. The absence of recognition may lead to dire consequences, for example, if, due to the lack of funded maintenance, the underlying libraries become vectors for malware attacks.
There have been proposals to assign credit to these packages using the dependency structure to calculate what has been termed transitive credit [16, 17] but so far, the problem has not yet been solved. To implement the proposed measure, software projects must start to publicize the packages they rely on in a way that enables recognition and citation, beyond the technical dependency already recorded in manifest files such as pyproject.toml, DESCRIPTION, cargo.toml, pom.xml, etc. One way for projects to do this is the inclusion of citation information for their own software outputs as well as for their direct dependencies, e.g., a citation file in the Citation File Format (CFF) [18]. Were such citation information available for the complete dependency stack of a program, transitive credit could be implemented by building weighted software citation networks [19].
@@
Following these ideas, we can put the software packages into a two-dimensional plane with the axes corresponding to the frequency of software mentions and a measure of network centrality (Figure 1). The majority of the packages will likely occupy the lower left corner of the plot, having a small number of mentions from the authors of scientific papers and limited centrality from usage in other software packages. Popular packages are used by many authors. We are interested in the “Nebraska” packages, not very visible, but of critical importance 111We use the term “Nebraska” packages throughout this article in reference to the original XKCD cartoon which noted that much of highly critical software is often less visible and less well known [15].. Of course, one can think about highly visible and critical “Pasteur” packages, which are used both directly and as a foundation for other libraries.
@@ 2.2 Network analysis
We additionally provide analysis for three variations of our network: unweighted, weighted, and the largest connected component from the weighted graph. The unweighted form of our network ignores the weights added to the edges between the seeding articles and their software mentions. The weighted network includes the article citation count weights. The difference between the unweighted and weighted versions of our analysis, allow us to parse different subtypes of the results. Specifically, comparing the weighted and unweighted networks allows to identify software which is only important due to its utilization and reference in high citation count articles.
@@ 3 Results
We found a dense core of popular packages that receive many mentions (e.g. ggplot2 in CRAN [33], tophat in PyPI and limma in Bioconductor [40, 41]), some of which have many dependencies themselves (e.g. ggplot2 [33]). We also found packages specific to certain communities (e.g. PRISMA in CRAN [35] or pymol in PyPI [37]).
@@
“Pasteur”: the highest sum of mentions and centrality percentile. To provide a “sanity check”, we added the data for three important packages: numpy, scipy, and napari.
@@
These metrics calculated for all packages are available as CSV files in the repository [22] (https://github.com/borisveytsman/SoftwareImpactHackathon2023_Tracing_dependencies/tree/main/data).
@@
| scipy | pypi | ||||
| numpy | pypi | ||||
| napari | pypi | ||||
@@
| scipy | pypi | ||||
| numpy | pypi | ||||
| napari | pypi | ||||
@@
| scipy | pypi | ||||
| numpy | pypi | ||||
| napari | pypi | ||||
@@ 4 Discussion
To check the results, we included in the tables three popular Python packages: scipy, numpy, and napari. We see that scipy and numpy are in high percentiles of package by centrality and mentions for all three versions of network studied. The newer and more specialized package napari is lower in the ratings, but still is above 50% percentile. All this seems reasonable.
Similarly we were pleased to note that the analysis identified the ‘tifffile‘ package as part of the Nebraska quadrant and anecdotes and quick inspection of the online materials confirm this as a high quality, widely used package with one person development team, albeit in California and not physically in Nebraska.
@@
The network analysis is limited by the quality of disambiguation and linking in the CZ Software Mentions dataset (see the discussion in [10]). Homonyms (different software packages with the same or close names) may change the network statistics, and preliminary research points to considerable potential for incorrect linking in the CZ Software Mentions dataset due to homonyms [6]. Notably, this affects the packages velvet and tophat (see Figure 2a), which have been falsely linked to Python packages. velvet was linked to a package that provides signal processing and communications algorithms in Python [36]. Meanwhile, all of the mentions of “velvet” point to different versions of algorithms for de novo short-read assembly using de Bruijn graphs in genomics, as described in [51] and implemented mostly in C [50]. tophat was linked to a package that provides a framework for collaborative and multiplayer mobile applications [52]. However, over 99 % of mentions point to different software called TopHat, a splice junction mapper for RNA-Seq reads, implemented mostly in C++[53, 54], and the remainder to the Top-Hat transform algorithm used for baseline subtraction in image processing. While the linking errors do not change the centrality values and our interpretation of them – both rely on disambiguation, not linking – they led to mislabelling of different, mostly non-PyPI packages as PyPI packages for the visualization in Figure 2.
@@
Our approach is limited to the open-source packages available in one of the chosen ecosystems. This excludes proprietary software like Excel or GraphPad Prism [55], and open-source non-package software like Gephi [56] (a complied application). In addition, our work excludes software projects such as BLAS [57], LAPACK [58] or the Linux Kernel itself. These packages are critically important but are often lower-level system dependencies that may not be clearly cited or referenced in an article’s text or provided in a software library’s dependency list. Without such references or dependencies, it is hard to detect their specific utilization. Without authors including references to lower-level software libraries in their articles, and without explicit dependency on lower-level libraries from mentioned software, we simply cannot include such software in our networks. The number of missing packages from the ecosystems investigated may be estimated from Table 14 of [10], which estimates the coverage for GitHub at 64.39%, the coverage for CRAN at 8.36%, the coverage for PyPI at 5.86%, and the coverage for Bioconductor at 3.23%. GitHub coverage can be used as a proxy for the share of the open source software: while some packages—perhaps growing in number—are not hosted by GitHub, their fraction remains low [59]. If we assume that each package exists only in one ecosystem (there are packages present in several, but we neglect this effect), the combined coverage by CRAN, PyPI, and Bioconductor is 17.45%, i.e., between a quarter and one third of the open source packages. It is very likely that each of these non-package pieces of software do depend on packages in our focus ecosystems. For open-source non-package software, it may be feasible to identify source repository URLs and then identify packages depended on by the software. That would need to be done by directly inspecting the source code, possibly using Software Bill of Materials (SBOM) tools [12]. On the other hand, some non-package software does not have source code available (including proprietary GUI software such as SPSS, and cloud-based services). Nonetheless, these pieces of code may very well be based on packages in our focal ecosystem (especially if the latter use non-restrictive licenses such as BSD, MIT). In the long term, there is a possibility that the requirement to provide SBOMs for U.S. federal government purchasers222as per Executive Order 14028: Improving the Nation’s Cybersecurity from May 12, 2021 will give insight into the packages on which proprietary software relies [60].
@@
Our current approach does not filter or subset the network into different time periods (i.e. by year) and as such, some of the software identified as central to our network such as velvet and tophat are perhaps historical artifacts that have not been updated in many years. These examples demonstrate that future analysis of software dependency networks should consider temporal data to understand how particular software projects (and more broadly, computational methods) rise in adoption and then decline as new methods become available.
10.
There are different software package repositories and ecosystems for Python and R with greatly overlapping dependency graphs. We used PyPI, CRAN and Bioconductor for this study. Alternative repositories that provide packages for Python and R, such as Conda channels (conda-forge, bioconda, etc.), are also worth investigating as there may be some software projects distributed solely through Conda channels. Conda provides alternate packaging standards from PyPI and CRAN and as such, over a large distribution, the same set of packages sourced from PyPI may have a slightly different dependency graph than one sourced from Conda. Conda also has more cross-language dependencies listed and may allow deeper traversal to find infrastructural libraries like BLAS; other improvements could include using PyPI Wheels or HomeBrew dependency systems.
@@ Acknowledgments
The code for this paper is available at: https://github.com/borisveytsman/SoftwareImpactHackathon2023_Tracing_dependencies [22], and the data are available at: https://zenodo.org/records/10048132 [25].